# BabyReasoningBench: Generating Developmentally-Inspired Reasoning Tasks for Evaluating Baby Language Models

### v1

@@ 2  BabyReasoningBench
False Belief Sally Anne 36.36 27.27
Counterfactual Possibilities 18.18 27.27
Transitive Inference 9.09 45.45
Physical Cause Effect 45.45 100.00
Control of Variables Strategy 18.18 27.27
Counterfactual Syllogism Pretend 9.09 63.64
Category Based Induction 45.45 63.64
Story Analogy Relational Shift 18.18 36.36
Causal Structure Learning 0.00 18.18
Blicket Detector Inference 0.00 63.64
Simple Causal Analogy 27.27 63.64
Simple Counterfactual Causal 9.09 63.64
Analogical Problem Solving 0.00 9.09
Class Inclusion Wording 0.00 0.00
Conservation of Number Accidental 100.00 90.91
Violation of Expectation False Belief 27.27 63.64
Exploratory Play Causal 36.36 18.18
False Belief Vignette Battery 54.55 45.45
False Belief Unexpected Transfer 0.00 9.09
@@ 3 Results
Table 1 and Figure 1 report task-wise accuracy for two BabyLM GPT-2 baselines. We find that the overall performance is low, but the pattern is strongly non-uniform across paradigms. The average accuracy increases from 28.2% for the 10M model to 40.7% for the 100M model, indicating that additional pre-training data compute help on this developmentally grounded suite.
Scaling primarily benefits causal and physical reasoning.
The largest gains from 10M100M appear in tasks that require tracking causal regularities across observations or applying simple physical principles. The 100M pretrained model reaches 100% on physical-cause-effect (vs. 45.45% at 10M) and improves substantially on evidence-integration tasks such as blicket-detector-inference (0.0063.64) and causal-structure-learning (0.0018.18). Similarly, simple-causal-analogy (27.2763.64) and simple-counterfactual-causal (9.0963.64) show large gains, suggesting that increased data compute supports more reliable use of local causal structure in short narratives.
Belief attribution remains difficult and inconsistent.
Across explicit false-belief paradigms, both models struggle. The 100M pretrained model improves on the implicit violation-of-expectation-false-belief items (27.2763.64), but explicit belief attribution remains weak: false-belief-unexpected-transfer stays near floor (0.009.09) and false-belief-sally-anne does not improve (36.3627.27). The mixed pattern mirrors the developmental literature motivating these paradigms: different task formats place different demands on language, inhibition, and representation of mental states, and success in one format does not guarantee robust belief reasoning across formats.
Pragmatics-sensitive tasks expose systematic failure modes.
The class-inclusion wording manipulation (class-inclusion-wording) is at 0.00% for both models. We feel this is consistent with the idea that these items are sensitive to pragmatic interpretation, world knowledge, and quantifier scope rather than pure set reasoning. Relatedly, performance on several reasoning primitives is modest and variable (counterfactual-possibilities, control-of-variables-strategy, story-analogy-relational-shift), suggesting that these models often fail to maintain task-relevant relational structure when it must be inferred rather than stated.
BabyReasoningBench reveals clear dissociations: scaling pre-training data helps in several causal and physical paradigms and some implicit-belief formats, while explicit false-belief and pragmatics-sensitive reasoning remain challenging. These patterns support the benchmark’s goal of going beyond a single “reasoning score” to diagnose which developmental-like competencies are (and are not) supported by child-directed training regimes.
@@ 4 Conclusion
We introduced  BabyReasoningBench, a benchmark of simple developmental reasoning tasks designed to evaluate baby language models inspired by classic experimental paradigms from developmental psychology. Since most of these tasks, at least framed in a textual fashion, are extremely easy for modern frontier models, we resort to them to create such a benchmark. The benchmark operationalizes theory of mind, analogical reasoning, causal inference, and several core reasoning primitives through multiple-choice questions using an LLM (GPT-5.2), enabling systematic variation of surface form while preserving experimental logic.
Baseline results on two BabyLM GPT-2 models show that (i) performance is substantially below ceiling overall, (ii) increases in pre-training data yield meaningful gains on multiple causal and physical reasoning tasks, but (iii) belief attribution and pragmatics-sensitive tasks remain difficult and sometimes near floor. The task-level dissociations—including sharp failures on class inclusion wording and inconsistent outcomes across false-belief formats—underscore why fine-grained developmentally grounded evaluation is valuable: it can separate reasoning competence from confounds tied to language form, framing, and inhibitory demands, and it can expose brittle generalization that would be invisible in aggregate benchmarks.

### v2

@@ 2  BabyReasoningBench
False Belief Sally Anne 63.64 63.64
Counterfactual Possibilities 63.64 54.55
Transitive Inference 9.09 36.36
Physical Cause Effect 63.64 100.00
Control of Variables Strategy 45.45 27.27
Counterfactual Syllogism Pretend 63.64 9.09
Category Based Induction 63.64 63.64
Story Analogy Relational Shift 27.27 45.45
Causal Structure Learning 81.82 18.18
Blicket Detector Inference 63.64 72.73
Simple Causal Analogy 54.55 72.73
Simple Counterfactual Causal 18.18 63.64
Analogical Problem Solving 0.00 36.36
Class Inclusion Wording 36.36 27.27
Conservation of Number Accidental 100.00 54.55
Violation of Expectation False Belief 81.82 90.91
Exploratory Play Causal 54.55 36.36
False Belief Vignette Battery 63.64 81.82
False Belief Unexpected Transfer 36.36 63.64
@@ 3 Results
Table 1 and Figure 1 report task-wise accuracy for two BabyLM GPT-2 baselines. We find moderate to strong overall performance for both models, with only a modest difference between scales. The average accuracy is 52.15% for the 10M model and 53.59% for the 100M model, indicating that both models solve a meaningful subset of the benchmark, but in a highly non-uniform way across paradigms.
Scaling yields selective gains, especially on some physical, causal, and analogy tasks.
The 100M model outperforms the 10M model on several tasks that involve local causal structure or relational transfer. It reaches 100.00% on physical-cause-effect (vs. 63.64% at 10M), improves on blicket-detector-inference (63.6472.73), simple-causal-analogy (54.5572.73), simple-counterfactual-causal (18.1863.64), and analogical-problem-solving (0.0036.36), and also performs strongly on violation-of-expectation-false-belief (90.91%). These gains suggest that additional pre-training can help on tasks that require short-range evidence integration or belief-consistent prediction.
At the same time, the benchmark does not show monotonic scaling.
The 10M model matches or exceeds the 100M model on a number of tasks, including counterfactual-possibilities, control-of-variables-strategy, counterfactual-syllogism-pretend, causal-structure-learning, class-inclusion-wording, conservation-of-number-accidental, and exploratory-play-causal. The largest reversal appears in causal-structure-learning, where the 10M model scores 81.82% and the 100M model only 18.18%. This pattern indicates that developmentally grounded reasoning performance is not well captured by a simple “larger is better” account, and may depend strongly on task framing and model-specific heuristics.
Belief attribution and pragmatics-sensitive reasoning are fragile rather than absent.
The rerun results also revise the interpretation of theory-of-mind-related tasks. Both models are now clearly above floor on explicit false-belief paradigms: false-belief-sally-anne is 63.64% for both models, false-belief-vignette-battery reaches 63.64% and 81.82%, and false-belief-unexpected-transfer reaches 36.36% and 63.64%. Likewise, class-inclusion-wording is no longer at floor, though performance remains modest. This suggests that these models can exhibit some belief-sensitive and pragmatics-sensitive behavior, but that such competence is unstable across formats rather than robustly generalized.
Overall,  BabyReasoningBench continues to reveal strong dissociations across developmental reasoning paradigms, but the updated results point to a more nuanced conclusion: both models show broader competence but overall still moderate performance, while scaling yields only selective and inconsistent benefits. This reinforces the value of fine-grained developmental evaluation for distinguishing robust reasoning abilities from brittle task-specific successes.
@@ 4 Conclusion
We introduced BabyReasoningBench, a benchmark of simple developmental reasoning tasks designed to evaluate baby language models inspired by classic experimental paradigms from developmental psychology. Because most of these tasks, when framed textually, are trivial for modern frontier models, they are especially useful as a diagnostic benchmark for small, developmentally trained models. The benchmark operationalizes theory of mind, analogical reasoning, causal inference, and several core reasoning primitives through multiple-choice questions, with task variants generated using an LLM (GPT-5.2) to systematically vary surface form while preserving logic.
Baseline results on two BabyLM GPT-2 models show that (i) performance is substantially below ceiling overall, (ii) increases in pre-training data yield meaningful gains on multiple causal and physical reasoning tasks, but (iii) belief attribution and pragmatics-sensitive tasks remain fragile and highly format-dependent rather than uniformly absent. The task-level dissociations—including sharp failures on class inclusion wording and inconsistent outcomes across false-belief formats—underscore why fine-grained developmentally grounded evaluation is valuable: it can separate reasoning competence from confounds tied to language form, framing, and inhibitory demands, and it can expose brittle generalization that would be invisible in aggregate benchmarks.